### v1

@@ 1. Introduction
The opinion dynamics is a research topic that has been quite active for several decades from various points of view: economical [6] [16], political [1] [13], mathematical [10] [2], sociological [17] [11],… Their study attempt to explain complex human behaviours when it comes to shaping individual and/or group opinions. To this end, different types of mathematical models [3] have been developed which, by their nature, can only represent very specific aspects of the mechanisms involved in the formation of opinion. Such models usually implement rules to describe the interactions between the agents involved, which may correspond to individuals, advertising messages, bots on social networks, or groups of any of these. In addition, the opinions of the agents can obey the general rules of each model or particular rules that only affect them, even if the rest of the agents interact with them following the general rules.
@@
In this paper we will study the influence of groups of manipulative agents on discrete-time bounded confidence models such as the Deffuant–Weisbuch [4] and Hegselmann–Krause [7] models as well as the Weighted Bounded Confidence models developed in [15], [12]. We will consider the number of manipulators and the rate of change of their opinions as the fundamental variables on which our study will be based. The models studied also have dynamics that lead to different types of final distributions: a single central cluster (consensus), multiple clusters distributed throughout the opinion space, two clusters at the extremes (polarization) or even oscillating opinions in a certain range of the opinion space. For this reason, there is no general definition of the Overton Window that can be suitable for all types of final distribution. Thus, we have chosen to measure only one central point (the mean) and dispersion (the standard deviation), although the latter value does not provide much information in the polarized distributions.
The article is organized as follows. Section 2 is devoted to introduce the definitions of the different types of agents as well as the models and their dynamics. Some simulations under different parameters are presented, which serve as examples of the already known dynamics under the influence of only normal agents or a group of stubborn agents. In section 3, the dynamics of the different models under the influence of manipulators are studied. Examples are provided to illustrate the different behaviours of opinions. These examples are used to interpret the variation of the statistical values of the final distribution of opinions obtained under the action of the manipulative group. In addition, for the Hegselmann–Krause model, the conditions that ensure the interaction between one normal group and another manipulative are established based on their number of agents and the speed at which the manipulators change their minds. Finally, in section 4, the results obtained and the work that can be carried out in future articles are discussed.
2. Models of opinion dynamics under bounded confidence
@@
In this section we will introduce the agent–based mathematical models of bounded confidence and the notation that we will use in the study of the opinion dynamics of the subsequent sections.
@@
Let us consider a discrete agent space $P=\{p_{i}\}_{i=1}^{N}$. Each agent $p_{i}\in P$ is a subject having an opinion about some issue. The opinions can be parametrized by some real number $x\in\mathbb{R}$ (that can be thought as the amount of resources, capital…, that the agent is willing to invest or increase from a budget), in such a way that $x(p_{i})=x_{i}$. We can consider that these values are all contained in an interval $[a,b]\subset\mathbb{R}$ called the opinion space, which can be interpreted as the range of all possible opinions. The opinion of each agent can vary along the discrete time $t=0,1,2,\ldots$, so we can denote by $x_{i}(t)$ the opinion of the agent $p_{i}$ at time $t$. The main feature of bounded confidence models, or BC–models for short (see [7]), is that each agent can only be influenced by agents with similar opinion. Indeed, for the agent $p_{i}\in P$ and for some value $\varepsilon_{i}>0$, called confidence threshold, we can denote the set of confidence of $p_{i}$ by $J_{i}(t)\in\mathbb{R}$ which can be defined by
@@
and it is made up of the indices of agents who have their opinion at a distance no higher than the confidence threshold $\varepsilon_{i}$ of the agent $p_{i}$. The number of agents in $J_{i}(t)$ will be denoted by #$\#\left(J_{i}(t)\right)$. Observe that, #$\#\left(J_{i}(t)\right)>0$ because $i\in J_{i}(t)$ for all $i=1,\ldots,N$. The dynamics of the opinion distribution $\{x_{i}(t)\}_{i=1}^{N}$ is described, for each $x_{i}$, by an iteration such that
@@
where the functions $F_{i}$ characterize the model defining the variation of the opinion of $p_{i}$ at the discrete time $t$ and where we have denoted $\mathbf{x}(t)=\left(x_{1}(t),\ldots,x_{N}(t)\right)\in\mathbb{R}^{N}$.
There are many ways to modify these models to adapt them to different cases. For example, the number of modified opinions at every iteration $t$ leads to different models, so we will say that a model is simultaneous if the iteration given in (2.2) is applied to all agents in order to modify their opinions for each $t$. On the other hand, we will call a model alternating if, for each $t$, iteration (2.2) is applied to a single agent. It is also possible to consider the existence of repulsive opinions, that is, an opinion $x$ is repulsive for the agent $p_{i}$ if $p_{i}$ modifies his opinion $x_{i}$ by moving away from $x$.
Another important modification in the models is setting up some types of agents which do not follow the dynamics of the rest of agents. In [8], [12], several types of agents are defined for the BC–model depending on where their opinions are situated in the opinion space, whether or not their opinions evolve under the dynamics equation (2.3) and their influence over the rest of agents. So, we can define the following types of agents:
  • A normal agent is that whose opinion evolves as $x_{i}$ under the dynamics equation (2.2). They set the general type of agents and influence each other in a homogeneous way.
  • A stubborn agent is that whose opinion do not change over time. The opinions of these agents influence others but are not influenced by the rest. So, if $p_{j}$ is a stubborn agent, then $x_{j}(t+1)=x_{j}(t)$ for all $t=0,1,2,\ldots$
  • A radical agent is a stubborn agent whose opinion is close to the boundary of the opinion space $[a,b]\subset\mathbb{R}$, that is, if $p_{j}$ is a radical agent, then $x_{j}(t)=c_{j}$ constant such that $c_{j}\approx a$ or $c_{j}\approx b$ for all $t\geq 0$.
  • A charismatic agent is a stubborn agent who is not radical.
If there are $1<K\in\mathbb{N}$ stubborn agents with the same opinion in the BC–model, they can be considered as a group of agents whose influence on the neighbouring agents is equivalent to the one of an only agent with the corresponding weighted opinion. So, radical and charismatic groups are treated in [8]. Such groups can be seen as an information campaign rather than a group of actors, so they influence society but are not counted as part of it. Notice that the distinction between radical and charismatics agents is not clearly objective because we have not quantified how close to the boundary of the opinion space $[a,b]$ must be an agent to be radical.
@@
  • A manipulative agent or manipulator is an agent $p_{i}$ whose opinion is not influenced by the rest of agents but follows a non–constant path $x_{i}(t)=f(t)\in[a,b]\in\mathbb{R}$, where $f$ is a function with values in the opinion space $[a,b]\in\mathbb{R}$.
@@
In this article, we will study some variants of the classic models of Deffuant–Weisbuch [4] and Hegselmann–Krause [7] as well as the models introduced by Toccaceli et al. [15]. The first two are simple and typical models based on an iteration such that
(2.3) $x_{i}(t+1)=\sum_{k\in\Omega_{i}(t)}w_{ik}x_{k}(t)$
corresponding to the weighted arithmetical mean value of the opinions of the agents in a subset $\Omega_{i}(t)\subset J_{i}(t)$ of the set of confidence of $x_{i}(t)$ where $w=(w_{ij})$ is a matrix such that $w_{ij}\in[0,1]$ and $\sum_{k}w_{ik}=1$. In this case, the choosing of the set $\Omega_{i}$ and the weights $w_{ij}$ characterize the model of opinion.
Throughout this article, we will consider the confidence thresholds of all agents to be equal, that is $\varepsilon_{i}=\varepsilon$ for all $i=1,\ldots,N$.
2.1. The Deffuant–Weisbuch model
According to equation (2.3), when for the agent $p_{i}$ we choose the set $\Omega_{i}=\{i,j\}$ where $j\in\{1,\ldots,n\}$ is a random index such that $j\neq i$ and the weight are $w_{ij}=1/2$, then we obtain the Deffuant–Weisbuch model [4], or DW–model for short, which explicitly is
This model can be implemented both in simultaneous and in alternating ways. Thus, at each iteration $t$ of the simultaneous model, a random agent $p_{j}$ is chosen for every agent $p_{i}$ with $i=1,\ldots,N$ and $i\neq j$, therefore every agent can modify its opinion simultaneously at every iteration. In the alternating model, at each iteration $t$, for a random agent $p_{i}$, another random agent $p_{j}$ is chosen to carry out the iteration (2.4). So, for each $t$, only one agent can modify its opinion. Obviously, these two ways of implementation are not equivalents although the opinion dynamics are similar. The main feature in which both forms of deployment differ is the speed of convergence. Since, in the second method, each agent intervenes in the iteration, then its convergence speed is much higher.
In figure 1, the opinion dynamics for a simultaneous Deffuant–Weisbuch model is shown with an equispaced initial distribution given by $x_{i}(0)=-1+\frac{2i}{101}$ for $i=1,\ldots,100$ and three different constant confidence thresholds $\varepsilon=0.1,0.45,0.6$. The distribution of opinions is considered stable whenever the values in each cluster differ less than $10^{-3}$ units achieved in $t_{max}$ iterations. Observe that the distances among final clustered opinions are greater than $\varepsilon$.
Figure 1. Example of opinion dynamics of the simultaneous Deffuant–Weisbuch model for a initial distribution of $N=100$ equispaced points $x_{i}(0)=-1+\frac{2i}{101}\in[-1,1]$ with $\varepsilon_{i}=\varepsilon\in\{0.1,0.45,0.6\}$ constant for all $i=1,\ldots,100$.
If a distribution of opinion interacts with a group of stubborn agents with constant opinion $x(t)=c\in[a,b]$ for all $t\geq 0$, then the opinion of agents in the confidence interval of the stubborn agents converges to this group’s opinion. A simulation of this situation can be seen in figure 2 for the same initial distribution as in figure 1 and a group of 20 stubborn agents (marked by a dashed line) with constant opinion $x(t)=0.75$ for all $t\geq 0$. Observe that the stubborn group attracts the neighbouring normal agents to its opinion.
Figure 2. Example of opinion dynamics of the simultaneous Deffuant–Weisbuch model for the same parameters as in figure 1 with a group of 20 stubborn agents with opinion $x=0.75$. The opinion of the stubborn group is drawn with a dashed line.
2.2. The Hegselmann–Krause model
In the case of the Hegselmann–Krause model [7], or HK–model, for the agent $p_{i}$, we consider the set $\Omega_{i}$ to be the whole confidence set $J_{i}$ and the new opinion $x_{i}$ of the agent $p_{i}$ is just the arithmetic mean values of all agents in $J_{i}$. So, we get
(2.5) $x_{i}(t+1)=\frac{1}{\#\left(J_{i}(t)\right)}\sum_{j\in J_{i}(t)}x_{j}(t)$
Again, analogously as for DW–model, the HK–model can be implemented as simultaneous or alternating.
To illustrate the opinion dynamics in this model, in the figure 3, we consider three simulations for an equispaced initial distribution $\{x_{i}(0)\}_{i=1}^{N}\subset[-1,1]$ for $N=100$ agents, with different constant confidence thresholds $\varepsilon\in\{0.1,0.45,0.6\}$ and $T=20$ iterations. Observe that for small values of $\varepsilon$ clusters appear at the final distribution $\{x_{i}(T)\}_{i=1}^{N}$ whereas for larger values of $\varepsilon$, the number of cluster decreases to 1, that is, consensus is reached. Note also that for intermediate values of $\varepsilon$, some clusters may appear that are not stable over time and end up grouping, reaching a consensus, after a relatively large number of iterations (see case $\varepsilon=0.45$ in the centre of figure 3), see [7].
Figure 3. Opinion dynamics in the simultaneous HK–model for a initial distribution of 100 equispaced points $x_{i}(0)=-1+\frac{2i}{101}$ for $i=1,\ldots,100$. The constant values of $\varepsilon$ are specified below every figure.
Figure 4. Opinion dynamics in the simultaneous HK–model for the same initial distribution as figure 3 with one group of 20 stubborn agents each with opinion at $x=0.75$ (marked with a dashed line).
2.3. The Weighted HK–models
It is possible to modify the BC–model assuming that the opinions change according to a weight function of the neighbouring opinions [5] in the form (2.2). In order to build an agent–based model to study the influence of fake news on society, Toccaceli et al. [15], Milli [12] define different models as in (2.2) that we will present below.
The Weighted Hegselmann–Krause model for group interactions is defined, in the opinion space $[-1,1]$, by the iteration
(2.6) $x_{i}(t+1)=\left\{\begin{array}[]{lcl}\displaystyle{x_{i}(t)+\frac{(1-|x_{i}(t)|)}{\#\left(\Omega_{i}(t)\right)}\sum_{j\in\Omega_{i}(t)}w_{ij}x_{j}(t)}&&\text{ if }\Omega_{i}(t)\neq\varnothing\\ x_{i}(t)&&\text{ otherwise }\end{array}\right.$
which modifies the value $x_{i}(t)$ by a weighted mean of the neighbouring opinions in $\Omega_{i}(t)=\{j:|x_{i}(t)-x_{j}(t)|\leq\varepsilon\text{ and }j\neq i\}$. The components of the matrix of weights $w=(w_{ij})$ describe the trustworthiness the agent $p_{i}$ attributes to the $p_{j}$’s opinion for $j\in\Omega_{i}(t)$ and their values will be taken as $w_{ij}\in[0,1]$.
Note that the term $\phi_{i}(t)=1-|x_{i}(t)|$ can be thought of as the probability density function $g(x)=1-|x|$ that gives weight to the interval $[-1,1]$. This makes moderate opinions (close to $x=0$) more likely to be modified and, therefore, after stabilization of opinion dynamics, most opinions are found in clusters $x=\pm 1$. Some opinions may fall outside these radical groups depending on the values of the confidence threshold $\varepsilon$ If the value of $\varepsilon$ is large enough, the final distribution is bimodal, polarized at $x=\pm 1$.
In figure 5, examples of opinion dynamics are shown for Weighted Hegselmann–Krause model for group interactions. The upper row shows a simulation for $N=100$ equispaced opinions of normal agents, confidence thresholds $\varepsilon=0.1,0.45,0.6$ and a random weight matrix $w=(w_{ij})$ but implemented in the alternating way (the opinion of a unique agent could be modified in each iteration). The lower row shows the simulation with the same parameters and weight matrix $w$ but implemented in a simultaneous way (the opinions of all agents could be modified in each iteration).
Figure 5. Opinion dynamics under the Weighted HK-model for group interactions for an equispaced initial distribution of $N=100$ agents for three values of $\varepsilon$ and a random matrix $w=(w_{ij})$. The graphs in the upper row have been implemented in a alternating way and the ones in the lower in a simultaneous way.
Note that the convergence rate for the simultaneous model is higher than for the alternating one.
Another model is the Weighted Hegselmann–Krause model for pairwise interactions. For the agent $p_{i}$, a random opinion $x_{j}(t)$ in the confidence interval is selected and the opinion of $p_{i}$ is modified by
(2.7) $x_{i}(t+1)=x_{i}(t)+\frac{x_{i}(t)+w_{ij}x_{j}(t)}{2}(1-|x_{i}(t)|)$
As a example of implementation of this model, in figure 6, simulations in the alternating (upper row) and simultaneous ways (lower row) for $N=100$ equispaced opinions of normal agents with three different confidence thresholds $\varepsilon=0.1,0.45,0.6$ and the same random weight matrix $w=(w_{ij})$ are shown.
Figure 6. Opinion dynamics under the Weighted HK model for pairwise interactions for an equispaced initial opinion distribution of $N=100$ normal agents for three values of $\varepsilon$ and a random matrix $w=(w_{ij})$. The upper row has been implemented in a alternating way and the lower one in a simultaneous way.
Also in [15], [12] there are some other models comprising mechanisms of attraction and/or repulsion. Let us consider an agent $p_{i}$ and a random opinion $x_{j}(t)$ in the distribution chosen such that $j\neq i$. The attractive mechanism modifies the opinion $x_{i}(t)$ by
where $s_{ij}(t)=sign(x_{i}(t)-x_{j}(t))$. In this case, if the opinion $x_{j}(t)$ is contained in the confidence interval of $x_{i}(t)$, then there is an attractive influence of $x_{j}(t)$ on the opinion $x_{i}(t)$, that is, the new opinion $x_{i}(t+1)$ moves closer to $x_{j}(t)$. We will call Attractive Weighted HK model to the iteration (2.8) corresponding to the attractive mechanism.
Figure 7. Opinion dynamics under the Attractive, Repulsive and Combined Weighted HK model for an equispaced initial opinion distribution of $N=100$ normal agents, a threshold of confidence of $\varepsilon=0.1$ and the same random matrix $w=(w_{ij})$.
In the figure 7 and in order to compare the different dynamics, simulations of 100 iterations for the Attractive, Repulsive and combined (Attractive-Repulsive) models are shown for the same equispaced initial opinion distribution of $N=100$ normal agents with a confidence threshold of $\varepsilon=0.1$ and a random weight matrix $w=(w_{ij})$. When the value of $\varepsilon$ is small, the final state for the Attractive model consists of three big clusters and some individual opinions in between. The largest cluster has the opinion of about 45 normal agents and it is located at $x=0$. There are two smaller clusters located at $x=\pm 1$ with about 24 normal agents each one. The opinions of the rest of normal agents are isolated in between those three clusters. For Repulsive and Attractive–Repulsive models, the central opinions oscillate with no convergence.
Figure 8. Opinion dynamics under the Attractive–Repulsive Weighted HK model for an equispaced initial opinion distribution of $N=100$ normal agents for three values of $\varepsilon$ and a random matrix $w=(w_{ij})$. The upper row has been implemented in a alternating way and the lower one in a simultaneous way.
3. Opinion dynamics with manipulative agents
@@ 3.1. The Overton window
The dynamics of the model will change the initial opinion distribution and thus also the measurements that characterize its properties: mean, median,… and their associated deviations. These characteristics can be shifted in the opinion space, just as dispersion can increase or decrease, depending on the behaviour of the possible stubborn or manipulative agents. So, these changes in statistical values are reminiscent of the political theory of the Overton window. This theory, enunciated by J. P. Overton, can be applied to some political subjects in which opinions can be ordered in a opinion space in some interval $[a,b]\in\mathbb{R}$, as we do in this article. It establishes a spectrum of classification of opinions in which only a small range is considered acceptable. Outside that acceptable range, opinions can be classified as radical and unthinkable. Government policies will always be in the range of acceptability, otherwise the government will lose voters or popularity. In this theory, it is the opinions within the Overton window that enable political action. Therefore, the shift and expansion or contraction of the Overton window (of acceptable opinions) can cause previously accepted opinions to become radical or unthinkable and vice versa, affecting the policies that the government will be able to carry out.
@@
In order to describe the Overton window we can choose two characteristics: a center point (mean or median) and a range or a dispersion value, but there is no canonical way to define the limits of acceptable opinions. In this article, we can use various statistical values that may be appropriate to define them and study their dynamics for different models with manipulative agents, although we will focus mainly on the mean and standard deviation.
3.2. Including manipulative agents
Consider an agent space $P=\{p_{i}\}_{i=1}^{N}$ with distribution of opinion at time $t$ given by $\{x_{i}(t)\}_{i=1}^{N}$ and $K\in\mathbb{N}$ manipulative agents with opinions $\{f_{j}(t)\}_{j=1}^{K}$ at time $t$. Now, the extended set of confidence $\overline{J}_{i}(t)$ of the normal agent $p_{i}$ is defined by

### v2

@@ 1. Introduction
The opinion dynamics is a research topic that has been quite active for several decades from various points of view: economical [7, 19], political [1, 16], mathematical [13, 2], sociological [20, 14] among others [11, 17]. Its study focuses on the phenomena that appear as a result of the interaction between the opinions of different agents. To this end, different types of mathematical models [4] have been developed which, by their nature, can only represent very specific aspects of the mechanisms involved in the formation of opinion. Such models usually implement rules to describe the interactions between the agents involved, which may correspond to individuals, advertising messages, bots on social networks, or groups of any of these. In addition, the opinions of the agents can obey the general rules of each model or particular rules that only affect them, even if the rest of the agents interact with them following the general rules.
@@
In this paper we will study the influence of groups of manipulative agents on discrete time bounded confidence models such as the Deffuant–Weisbuch [5] and Hegselmann–Krause [9] models as well as the weighted bounded confidence models developed in [18, 15]. The models studied also have dynamics that lead to different types of final distributions: a single central cluster (consensus), multiple clusters distributed throughout the opinion space, two clusters at the extremes (polarization) or even oscillating opinions in a certain range of the opinion space. For this reason, there is no general definition of the Overton Window that can be suitable for all types of final distribution.
The article is organized as follows. Section 2 is devoted to introduce the motivations in which this paper is based and a general model with manipulative agents. The following sections, which are largely independent of each other, deal with the various models that we will study in the article. In section 3, we will describe the Hegselmann–Krause model with and without manipulative agents. The interaction between manipulative and normal agents is mathematically analysed and a simulation showing the ability of manipulators to modify normal opinions is presented. The study of this interaction is based on the number of manipulative agents and the speed at which they change their opinion. In a similar way, in section 4, the Deffuant–Weisbuch model is described and the influence of manipulators is simulated under analogue parameters. Section 5 describes three types of Weighted HK models: Attractive, Repulsive and Attractive–Repulsive. Again, the influence of one manipulative group is studied by simulation. Finally, in section 6, the results obtained and the work that can be carried out in future articles are discussed.
2. Bounded confidence models with manipulative agents
@@
The aim of this section is to introduce the general opinion dynamics model with manipulative agents as a modification of already known bounded confidence agent–based models.
2.1. Opinion dynamics without manipulative agents
@@
Let us consider a discrete agent space $P=\{p_{i}\}_{i=1}^{N}$. Each agent $p_{i}\in P$ is a subject having an opinion about some issue. The opinions can be parametrized by some real number $x\in\mathbb{R}$ (that can be thought as the amount of resources, capital, etc. that the agent is willing to invest or increase from a budget), in such a way that $x(p_{i})=x_{i}$. We can consider that these values are all contained in an interval $[a,b]\subset\mathbb{R}$ called the opinion space, which can be interpreted as the range of all possible opinions. The opinion of each agent can vary along the discrete time $t\in\mathbb{N}\cup\{0\}$, so we can denote by $x_{i}(t)$ the opinion of the agent $p_{i}$ at time $t$ and $\mathbf{x}(t)=\left(x_{1}(t),\ldots,x_{N}(t)\right)\in\mathbb{R}^{N}$. The main feature of bounded confidence models, or BC–models for short (see [9]), is that each agent can only be influenced by agents with close opinion. Indeed, for the agent $p_{i}\in P$ and for some value $\varepsilon_{i}>0$, called confidence threshold, we can denote the set of confidence of $p_{i}$ by $J_{i}(t)\in\mathbb{R}$ which can be defined by
@@
and it is made up of the indexes of agents who have their opinion at a distance no higher than the confidence threshold $\varepsilon_{i}$ of the agent $p_{i}$. The number of agents in $J_{i}(t)$ will be denoted by $|J_{i}(t)|$. Observe that, $|J_{i}(t)|>0$ because $i\in J_{i}(t)$ for all $i=1,\ldots,N$. The dynamics of the opinion distribution $\{x_{i}(t)\}_{i=1}^{N}$ is described, for each $x_{i}$, by an iteration such that
@@
where the functions $F_{i}$ depend only on the opinions $x_{j}(t)$ where $j\in\Omega_{i}(t)\subset J_{i}(t)$. The set $\Omega_{i}(t)$ is defined as the set of indexes of agents which influence the opinion of agent $p_{i}$. So, the opinion $x_{i}(t+1)$ is defined by the iteration (2.2) only if $j\in\Omega_{i}(t)$, otherwise $x_{i}(t+1)=x_{i}(t)$. This implies that the sets $\Omega_{i}$ and the functions $F_{i}$ characterize the model defining the variation of the opinion of $p_{i}$ at the discrete time $t$.
In addition to the choice of the sets $\Omega_{i}$ and the functions $F_{i}$, there are many ways to modify these models to adapt them to different cases. For example, the number of modified opinions at every iteration $t$ leads to different models, so we will say that a model is simultaneous if the iteration given in (2.2) is applied to all agents in order to modify their opinions for each $t$. On the other hand, we will call a model alternating if, for each $t$, iteration (2.2) is applied to a single random agent.
2.2. Types of agent
In [10, 15, 3], several types of agents are defined for the BC–model depending on where their opinions are situated in the opinion space, whether or not their opinions evolve under the dynamics equation (2.2) and their influence over the rest of agents. In this article, we will consider the following types of agents:
  • A normal agent is that whose opinion evolves as $x_{i}$ under the dynamics equation (2.2). They establish the general type of agents and influence each other homogeneously.
  • A stubborn agent is that whose opinion do not change over time. The opinions of these agents influence others but are not influenced by the rest. So, if $p_{j}$ is a stubborn agent, then $x_{j}(t+1)=x_{j}(t)$ for all $t\in\mathbb{N}\cup\{0\}$
If there are $1<K\in\mathbb{N}$ stubborn agents with the same opinion in the BC–model, they can be considered as a group of agents whose influence on the neighbouring agents is equivalent to the one of an only agent with the corresponding weighted opinion. Such groups can be seen as an information campaign rather than a group of actors, so they influence society but are not counted as part of it.
@@ 2.2. Types of agent
  • A manipulative agent or manipulator is an agent $p_{i}$ whose opinion is not influenced by the rest of agents but follows a non–constant path $x_{i}(t)=f_{i}(t)\in[a,b]\in\mathbb{R}$, where $f_{i}$ is a function with values in the opinion space $[a,b]\in\mathbb{R}$ that will be defined prior to the implementation of the model.
@@ 2.3. The Overton window
In the 1990s, J. P. Overton introduced a political theory, later called the Overton window, according to which government policies can be classified within a spectrum that describes their acceptability by society. It establishes a classification of opinions in which only a small range is considered acceptable. Outside that acceptable range, opinions can be classified as radical or even unthinkable. Government policies will always be in the range of acceptability, otherwise the government will lose voters or popularity. In this theory, it is the opinions within the Overton window that enable political action. Therefore, the shift and expansion or contraction of the Overton window (of acceptable opinions) can cause previously accepted opinions to become radical or unthinkable and vice versa, affecting the policies that the government will be able to carry out.
In relation to opinion models, we can think of the Overton window of acceptable opinions as a interval $[c,d]$ contained in the opinion space $I=[a,b]$ of all possible opinions, that is $[c,d]\subset I$. This subinterval can be characterized by two numbers: a center value and a width. But notice that there is no canonical way to define the limits of acceptable opinions and it also depends on the existence of a number of agents who have the same or similar opinion.
The dynamics of the model, at every iteration, will change the initial opinion distribution and thus also the measurements that characterize some properties: mean value, the position of possible final clusters as well as their number of agents contained in them. These characteristics can be shifted in the opinion space, just as dispersion can increase or decrease, depending on the behaviour of the possible stubborn or manipulative agents. So, these changes in statistical values are reminiscent of the movement of the Overton window in the political theory.
@@
In section 3.1, after introducing the Hegselmann–Krause model with manipulative agents, we will illustrate how a group of manipulators is able to influence the opinion of society to shift the range of opinions towards one of the extremes of the opinion space, even to move many opinions to the most radical possible opinion (see figure 3).
2.4. A general model with manipulative agents
According to the definition of manipulative agent in section 2.2, the opinion of the manipulators is not under the dynamics of the models, but follows a trajectory previously defined for each time $t$. Therefore, the manipulators can influence the normal agents, but the normal agents will not be able to influence the manipulators.
So, we will introduce manipulative agents in a general model. Let us fix the opinion space to $I=[-1,1]$. Consider an agent space $P=\{p_{i}\}_{i=1}^{N}$ with distribution of opinion at time $t$ given by $\{x_{i}(t)\}_{i=1}^{N}$ and $K\in\mathbb{N}$ manipulative agents with predefined opinions $\{f_{j}(t)\}_{j=1}^{K}$ for time $t$. Now, the extended set of confidence $\overline{J}_{i}(t)$ of the normal agent $p_{i}$ is defined by